# On the metric property of quantum Wasserstein divergences

### v3

@@ 4.1. Triangle inequality for qubits — the case of the symmetric cost
$D_{s}^{2}(\rho,\rho)=2\left(1-\sqrt{1-\left|\mathbf{b}_{\rho}\right|_{2}^{2}}\right)$ (35)
@@
$C_{s}=4\left(||\sigma_{1}\rangle\rangle\langle\langle\sigma_{1}||+||\sigma_{2}\rangle\rangle\langle\langle\sigma_{2}||+||\sigma_{3}\rangle\rangle\langle\langle\sigma_{3}||\right)$
@@
$\langle\langle\sqrt{\rho}||C_{s}||\sqrt{\rho}\rangle\rangle=4\left|\left|\frac{1}{2}(\sqrt{\lambda}-\sqrt{1-\lambda})\frac{\mathbf{b}_{\rho}}{\left|\mathbf{b}_{\rho}\right|_{2}}\cdot\mathbb{\sigma}\right|\right|_{HS}^{2}=$
$=\left(1-2\sqrt{\frac{1}{2}(1+\left|\mathbf{b}_{\rho}\right|_{2})\frac{1}{2}(1-\left|\mathbf{b}_{\rho}\right|_{2})}\right)\frac{2\left|\mathbf{b}_{\rho}\right|_{2}^{2}}{\left|\mathbf{b}_{\rho}\right|_{2}^{2}}=2\left(1-\sqrt{1-\left|\mathbf{b}_{\rho}\right|_{2}^{2}}\right).$
@@
$\leq 6-2\mathbf{b}_{\rho}\cdot\mathbf{b}_{\tau}-4\left|\mathbf{b}_{\rho}-\mathbf{b}_{\omega}\right|_{2}-4\left|\mathbf{b}_{\omega}-\mathbf{b}_{\tau}\right|_{2}+2\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)$
@@
$\geq 2\left(4\left|\mathbf{b}_{\rho}-\mathbf{b}_{\omega}\right|_{2}-\left(1-\sqrt{1-\left|\mathbf{b}_{\rho}\right|_{2}^{2}}\right)-\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)\right)^{\frac{1}{2}}\times$
×$\times\left(4\left|\mathbf{b}_{\omega}-\mathbf{b}_{\tau}\right|_{2}-\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)-\left(1-\sqrt{1-\left|\mathbf{b}_{\tau}\right|_{2}^{2}}\right)\right)^{\frac{1}{2}}.$
@@
$6-2\mathbf{b}_{\rho}\cdot\mathbf{b}_{\tau}-4\left|\mathbf{b}_{\rho}-\mathbf{b}_{\omega}\right|_{2}-4\left|\mathbf{b}_{\omega}-\mathbf{b}_{\tau}\right|_{2}+2\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)\leq$
$\leq 2\left(4\left|\mathbf{b}_{\rho}-\mathbf{b}_{\omega}\right|_{2}-\left(1-\sqrt{1-\left|\mathbf{b}_{\rho}\right|_{2}^{2}}\right)-\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)\right)^{\frac{1}{2}}\times$
×$\times\left(4\left|\mathbf{b}_{\omega}-\mathbf{b}_{\tau}\right|_{2}-\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)-\left(1-\sqrt{1-\left|\mathbf{b}_{\tau}\right|_{2}^{2}}\right)\right)^{\frac{1}{2}}.$ (36)
@@
There are various ways to obtain easy-to-check examples of states $\rho,\omega,\tau,\in\mathcal{S}(\mathbb{C}^{2})$ that satisfy the assumption of Corollary 4. One way is to fix $\omega=\frac{1}{2}I,$ that is, $\mathbf{b}_{\omega}=0.$ In this case, simple $2$-variable calculus shows that (36) is satisfied whenever both $\left|\mathbf{b}_{\rho}\right|_{2}$ and $\left|\mathbf{b}_{\tau}\right|_{2}$ are at least $\frac{1}{2},$ no matter what the angle between $\mathbf{b}_{\rho}$ and $\mathbf{b}_{\tau}$ is. Moreover, a numerical test shows if we choose a random triplet of states (the states are chosen independently according to the uniform distribution on the Bloch ball), it will satisfy (36) with high probability (around $96\%$). The details of this numerics is presented in the subfolder ”Corollary-4” of [8]. Note, however, that there are states that do not satisfy (36): the easiest example is $\rho=\omega=\tau=\frac{1}{2}I.$

### v4

@@ 4.1. Triangle inequality for qubits — the case of the symmetric cost
$D_{s}^{2}(\rho,\rho)=4\left(1-\sqrt{1-\left|\mathbf{b}_{\rho}\right|_{2}^{2}}\right)$ (35)
@@
$C_{s}=8\left(\frac{1}{2}||\sigma_{1}\rangle\rangle\langle\langle\sigma_{1}||+\frac{1}{2}||\sigma_{2}\rangle\rangle\langle\langle\sigma_{2}||+\frac{1}{2}||\sigma_{3}\rangle\rangle\langle\langle\sigma_{3}||\right)$
@@
$\langle\langle\sqrt{\rho}||C_{s}||\sqrt{\rho}\rangle\rangle=8\left|\left|\frac{1}{2}(\sqrt{\lambda}-\sqrt{1-\lambda})\frac{\mathbf{b}_{\rho}}{\left|\mathbf{b}_{\rho}\right|_{2}}\cdot\mathbb{\sigma}\right|\right|_{HS}^{2}=$
$=2\left(1-2\sqrt{\frac{1}{2}(1+\left|\mathbf{b}_{\rho}\right|_{2})\frac{1}{2}(1-\left|\mathbf{b}_{\rho}\right|_{2})}\right)\frac{2\left|\mathbf{b}_{\rho}\right|_{2}^{2}}{\left|\mathbf{b}_{\rho}\right|_{2}^{2}}=4\left(1-\sqrt{1-\left|\mathbf{b}_{\rho}\right|_{2}^{2}}\right).$
@@
$\leq 6-2\mathbf{b}_{\rho}\cdot\mathbf{b}_{\tau}-4\left|\mathbf{b}_{\rho}-\mathbf{b}_{\omega}\right|_{2}-4\left|\mathbf{b}_{\omega}-\mathbf{b}_{\tau}\right|_{2}+4\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)$
@@
$\geq 2\left(4\left|\mathbf{b}_{\rho}-\mathbf{b}_{\omega}\right|_{2}-2\left(1-\sqrt{1-\left|\mathbf{b}_{\rho}\right|_{2}^{2}}\right)-2\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)\right)^{\frac{1}{2}}\times$
×$\times\left(4\left|\mathbf{b}_{\omega}-\mathbf{b}_{\tau}\right|_{2}-2\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)-2\left(1-\sqrt{1-\left|\mathbf{b}_{\tau}\right|_{2}^{2}}\right)\right)^{\frac{1}{2}}.$
@@
$6-2\mathbf{b}_{\rho}\cdot\mathbf{b}_{\tau}-4\left|\mathbf{b}_{\rho}-\mathbf{b}_{\omega}\right|_{2}-4\left|\mathbf{b}_{\omega}-\mathbf{b}_{\tau}\right|_{2}+4\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)\leq$
$\leq 2\left(4\left|\mathbf{b}_{\rho}-\mathbf{b}_{\omega}\right|_{2}-2\left(1-\sqrt{1-\left|\mathbf{b}_{\rho}\right|_{2}^{2}}\right)-2\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)\right)^{\frac{1}{2}}\times$
×$\times\left(4\left|\mathbf{b}_{\omega}-\mathbf{b}_{\tau}\right|_{2}-2\left(1-\sqrt{1-\left|\mathbf{b}_{\omega}\right|_{2}^{2}}\right)-2\left(1-\sqrt{1-\left|\mathbf{b}_{\tau}\right|_{2}^{2}}\right)\right)^{\frac{1}{2}}.$ (36)
@@
There are various ways to obtain easy-to-check examples of states $\rho,\omega,\tau,\in\mathcal{S}(\mathbb{C}^{2})$ that satisfy the assumption of Corollary 4. One way is to fix $\omega=\frac{1}{2}I,$ that is, $\mathbf{b}_{\omega}=0.$ In this case, simple $2$-variable calculus shows that (36) is satisfied whenever both $\left|\mathbf{b}_{\rho}\right|_{2}$ and $\left|\mathbf{b}_{\tau}\right|_{2}$ are at least $\frac{1}{2},$ no matter what the angle between $\mathbf{b}_{\rho}$ and $\mathbf{b}_{\tau}$ is. Moreover, a numerical test shows if we choose a random triplet of states (the states are chosen independently according to the uniform distribution on the Bloch ball), it will satisfy (36) with high probability (around $85\%$). The details of this numerics is presented in the subfolder ”Corollary-4” of [8]. Note, however, that there are states that do not satisfy (36): the easiest example is $\rho=\omega=\tau=\frac{1}{2}I.$